OpenStudy (anonymous):

hello, how do you get the answer which is 1 3/11 when i came up with 1 3/22?? the question i have is 1 3/4 (8/11).

OpenStudy (anonymous):

i dont understand ur question

OpenStudy (anonymous):

the question says in the book: 1 3/4 (8/11) however the answer listed there says: 1 3/11 but i've come up with 1 3/22. wonder how to do a solution to get that answer in the book?

OpenStudy (anonymous):

1 3/4 = 7/4 7/4 (8/11) = 14/11= 1 3/11

OpenStudy (anonymous):

thank you but i still don't get how it comes to that answer, i mean where does 14 come int this, i've gathered 7/4 but after that i still don't get it. can you please explain even i try my best to understand? it can be a drawing if possible.

OpenStudy (anonymous):

7/4 x 8/ 11 = 56/44 ----> (56:4)/(44:4) = 14/11

OpenStudy (anonymous):

thank you.